News by Date
News by Category
Solar System
White Dwarfs
Neutron Stars
Black Holes
Milky Way Galaxy
Normal Galaxies
Galaxy Clusters
Cosmology/Deep Field
Press Resources
Status Reports
Press Advisories
Image Releases
Release Guidelines
Image Use Policy
Web Shortcuts
Chandra Blog
RSS Feed
Email Newsletter
News & Noteworthy
Image Use Policy
Questions & Answers
Glossary of Terms
Download Guide
Get Adobe Reader
Chandra @ NASA
Visit the Chandra pages at the NASA portal (opens in new window)
Image Use
Image Use Policy & Request Form
Guidelines for utilizing images, applets, movies, and animations featured in this Web Site.
Operations CXO Status Report

Friday 31 October 2003 7.00am EST

The observing schedule was significantly impacted last week due to the extremely high levels of solar activity. Following the manual execution of SCS107 on Oct 24, science was resumed on Oct 25 at 1:55pm EST but was halted by an autonomous execution of SCS107 on Oct 26 at 2:22pm EST. Loads were resumed but halted once more on Oct 28 at 9:02am EST. Science observing has not resumed since then due to continued high radiation levels. During the week two maneuver loads were uplinked to ensure proper attitude during perigee passage and perform momentum management. The team is monitoring the solar activity closely and has been developing a series of mission plans to ensure a rapid resumption of science when possible.

A Chandra image release was issued on Oct 28 that describes a series of observations of the spiral galaxy NGC 1637 that show the variable nature of the X-ray sources within the galaxy. For details see:

Of note last week was the Chandra Calibration Workshop and review held Oct 27-28. The workshop provided the community with the state of the art of the Chandra team's knowledge of the detectors and mirrors, and also encouraged the general community to participate in the process of calibrating the satellite (

The observing schedule for next week is shown below (assuming solar levels have subsided).

FS Tau AB                ACIS-S       Nov 2
IRAS FSC 01562+2528      ACIS-S       Nov 3
M31-Center               ACIS-I
HER X-1                  ACIS-S/HETG
G21.5-0.9                ACIS-I
G21.5-0.9                ACIS-S
G21.5-0.9                HRC-I
G21.5-0.9                HRC-S        Nov 4
alpha CrA                ACIS-I
SDSS J000230.71+004959.1 ACIS-S
Radiation Belts
PSR J1809-1917           ACIS-S       Nov 5
LBQS 0004+0147           ACIS-S
RCrADarkCloud            ACIS-I
SDSS J173551.92+535515.6 ACIS-S
0234+285                 ACIS-S       Nov 6
IRAS FSC 02524+2046      ACIS-S
MS0735.6+7421            ACIS-S
SDSS J171535.96+632336.0 ACIS-S
PC 1450+3404             ACIS-S
NGC4350                  ACIS-S       Nov 7
NGC 4216                 ACIS-S
Radiation Belts
LBQS 2358+0216           ACIS-S
HH34 IRS                 ACIS-S       Nov 8
0540-697                 ACIS-S
MU LEP                   HRC-S/LETG


All spacecraft subsystems continued to support nominal operations.

Return to Status Reports